A previous random sample of 4000 U.S. citizens yielded 2250 who are in favor of gun control legislation. How many citizens would need to be sampled for a 95% confidence interval to estimate the true proportion within 2%
Answer #1

Solution :

Given that,

n = 4000

x = 2250

\hat p = 2250 / 4000 = 0.5625

1 - \hat p = 1 - 0.5625 = 0.4375

margin of error = E = 2% = 0.02

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p)

= (1.96 / 0.02)2 * 0.5625 * 0.4375

= 2363.48

= 2364

2364 citizens would need to be sampled .